A finite math class of 50 students is composed of freshmen and sophomores according to the table below. A student from the class is selected at random.

Freshmen Sophomores
Females 12 8
Males 14 16

a) What is the probability that the selected student is either a male or a freshman?

b) What is the probability that the selected student is neither a female nor a sophomore?
Answer #1
data Freshmen Sophomores Total
Females 12 8 20
Males 14 16 30
Total 26.00 24.00 50

(a) P(either male or freshman) = P(male) + P(freshman)- P(male & freshman)

P(either male or freshman) = 30/ 50 + 26/50 - 14/50 = 42 / 50 = 0.840

P(either male or freshman) = 0.840

(b) P(neither femle nor a sophomore) = 14/ 50 = 0.28

  P(neither femle nor a sophomore) = 0.28